THE SIDE OF AN EQUILATERAL TRIANGLE ARE GIVEN AS(3x+2)cm, (2y-x)cm and (y+3)cm.find x and y, and the perimeter of the triangle.
Answer by Boreal(15235)

sides are equal
so 3x+2=2y-x=y+3
2y-x=y+3
y-3=x
substitute
3(y-3)+2=y+3
3y-9+2=y+3
2y=10
y=5
x=2
the sides are 8, 8, 8 with perimeter 24.